OpenStudy (anonymous):

Solve the equation. log(x + 7) – logx = 3

OpenStudy (anonymous):

@satellite73 could you please help?

OpenStudy (anonymous):

\[log(x + 7) – logx = 3 \] \[log[ \frac{(x + 7)}{x} ] = 3 \] \[\frac{x+7}{x} = 10^3 \] Convert to exponential form: \[\frac{x+7}{x} = 1000 \] \[x+7 = 1000x\] Solve for x

OpenStudy (anonymous):

is it .0070???

OpenStudy (anonymous):

x = 0.0070070

OpenStudy (anonymous):

So yes.

OpenStudy (anonymous):

thankyou so much!! could you help with another one please?

OpenStudy (anonymous):

I am not that good with logarithms I just happen to know to use the rule: \[\log_{b}\left(\frac{x}{y}\right) = \log_{b}(x) - \log_{b}(y)\]

OpenStudy (anonymous):

alright! well thank you!

OpenStudy (anonymous):

Np :)